OpenStudy (anonymous):

5. Air is being pumped into a balloon, at a rate represented by the equation y = t1/3 where y is the rate of airflow into the balloon in cm3/s and t is the time in seconds. The balloon already has 10 cm3 of air in it before air is pumped into it. The balloon will burst if there is more than 100cm3 of air in it. Find the time taken for the balloon to burst.

OpenStudy (anonymous):

is it y = t/3?

OpenStudy (anonymous):

90 = t/3 t = 270 seconds